OpenStudy (anonymous):

What are the steps to solve this problem? 0.000027 ÷ 0.000009 Prepare your response to this question. When you are ready, you may submit your response as an audio recording or as a written essay. Be sure to include the following: Requirements: • Explain how to use scientific notation to solve the problem. • Describe how to divide numbers written in scientific notation. • Give you final answer in simplest form. (5 points) Need help don't know how to do this at all, when they ask for steps I'm lost.

OpenStudy (anonymous):

I'm typing it if anyone can just help me with the steps to the problem, it's kinda hard cause I never used this math in life and I usually used a calculator anyways.

OpenStudy (anonymous):



OpenStudy (anonymous):

I understand that I just don't know how to explain the steps to someone

OpenStudy (anonymous):

I don't follow all the steps in my mind I just remember the important things in solving is what im saying, for example 2 + a = 4 I just know you subtract 2 to 4 to find a not the add by both sides etc. the details

OpenStudy (anonymous):

every little detail in these problems I forgot and they are long gone.

OpenStudy (anonymous):

okay so here is what you will say to someone.... The number of digits in numerator after decimal can be multiplied in 10s' in denominator and vice versa.... So numerator has 6 units after decimal and denominator has 6 units after decimal... You multiply that number of zeros in denominator and numerator simultaneousily so the final no is (0.000027x1000000 ÷ (0.000009x 1000000)= 27÷9 =3

OpenStudy (ranga):

What the teacher is expecting is for you to put the numerator and the denominator in scientific notation first and then do the division.